How much heat is necessary to change 10 g of ice at - 20 degrees * C into water at 10 degrees * C ? a. 0.02 kcal b . 0.2 kcal c. 1 kcal d. 0.08 kcal
Answer #1

Answer

Q1= energy needed to warm the ice from -20 degree C to 0 degree C

Q1= mass of ice*specific heat of ice *temperature change

    = (10 g)*(2.1 J/gC)*(20 degree C)= 420 J

Q2= energy needed to phase change of H2O(from solid to liquid )

Q2= mass of H2O* latent heat of fusion of water

      = (10 g)*(334 J/g)= 3340 J

Q3= energy needed to warm the water from 0 to 10 degree C

Q3= mass of water *specific heat of water *temperature change

     = (10 g)*(4.184 J/gC)*(10)= 418.4 J

Q= Q1+Q2+Q3 = 420+3340+418.4 = 4178.4 J

1 kcal = 4184 J

The answer is 1 Kcal option (C)